# MRIR: Integrating Multimodal Insights for Diffusion-based Realistic Image Restoration

### v1

@@ III-A Motivation and Overview
Figure 1: The comparison of different prompts and their corresponding restoration results with PASD [30].(a) -(d) show the null prompt, BLIP prompt from LQ image, BLIP prompt from GT image and MLLM prompt predicted from LQ image and its corresponding restoration result.
Overview. Based on the aforementioned observations, we propose a multimodal-perception approach for realistic image restoration. To fully leverage the generative priors of pre-trained T2I models such as stable diffusion, we utilize textual level and visual/pixel level information to guide the diffusion process in image restoration, aiming to generate realistic restoration results with coherent structures and clear details. The overall framework of our model is depicted in Fig 2. We employ stable diffusion as the backbone and control the diffusion process from both the textual level and pixel level. The textual level includes two branches: one branch utilizes the inference capability of a pre-trained multimodal language model to predict potential restoration results from the LQ images, generating restored text embeddings. Additionally, to ensure the semantic fidelity of the generated predictions to the original LQ images, we design an image embedding branch. An image encoder extracts image embedding, and a Refine Layer is employed to correct the image embedding and mitigate the influence of degraded information in LQ images. The text embedding and image embedding collaborate as the textual level information to control the restoration process of the diffusion model. The pixel level control consists of a Pixel-level Processor and a ControlNet module. The Pixel-level Processor extracts potential clear features from the low-quality image, while the ControlNet module utilizes these features to generate pixel control in the latent domain, thus enabling spatial structural control in the diffusion process. Finally, the denoising U-Net utilizes text embedding, image embedding, and pixel control to generate high-quality restoration results. Our model is trained end-to-end without requiring any additional training while SeeSR and CoSeR need additional training to obtain the semantic information from LQ images. With the aid of the powerful image understanding and reasoning capabilities of MLLM, our method can directly extract meaningful semantic information from LQ images without the need for additional training.
Figure 2: The framework of the proposed method, which consists of textual level control module, pixel level control module and the denoising U-Net.
III-B Textual Level Restoration with MLLM
Prompt is an important factor influencing the quality and details of the generation from stable diffusion. Directly extracting prompts from low-quality images using BLIP [53] can be affected by degradation artifacts, resulting in inaccurate representations. Therefore, extracting high-quality textual representations from the original low-quality images is an important issue to consider. Previous methods, such as SeeSR and CoSeR, employed additional training to distill semantic or cognitive information from LQ images. However, these methods require extra training and their effectiveness relies on the training dataset, limiting their generalization capability.
@@ III-B Textual Level Restoration with MLLM
Benefiting from multimodal large language models, which possess extensive knowledge and powerful perception and reasoning capabilities, we utilize an MLLM to predict reasonable textual representations for potential blind image restoration results. Large language models can generate promising responses to user queries based on image information. We employ the pre-trained multi-modal large language model LLaVA-7B  [^1] https://huggingface.co/liuhaotian/llava-v1.5-7b to predict potential clear representations from low-quality images, generating coherent high-quality textual descriptions and encoding them as text embeddings. The aforementioned process utilizes pre-trained models and does not require fine-tuning. The process can be summarized as follows:
@@
where $T_{text}$ and $I_{LQ}$ denote the text embedding and the input low-quality image respectively, $T_{input}$ denotes the text instruction to the MLLM, such as “Describe the image in a very detailed manner if we remove the degradation artifacts from the image.” and $E_{text}$ and $E_{MLLM}$ indicates the CLIP text encoder and MLLM model respectively. MLLM can take both image and text sequences as input and generate coherent descriptions as output. Compared to predicting semantic tags [31] and cognitive embeddings [32], the predicted descriptions are more detailed and conducive to human feedback.
The inference results of the MLLM include fine-grained descriptions of the restored image. To ensure that the generated image remains semantically faithful to the original low-quality image, we introduce an image embedding branch to extract visual information as a supplement. Specifically, we use CLIP image encoder to extract additional image embedding. To eliminate information about degradation artifacts present in the image embedding, we design a Refine Layer to correct the image embedding, generating image embedding that is semantically faithful to the original low-quality image but free from degradation information. The process can be expressed as follows:
@@
where $T_{image}$, $E_{image}$, and $E_{refine}$ represent the image embedding, CLIP image encoder, and the Refine Layer respectively. The refine layer contains three MLP layers with Layer Normalization and LeakyReLU activations (except for the final layer).
Ultimately, the inferred text embedding ($T_{text}$) and image embedding ($T_{image}$) collaborate as the textual level information to control the restoration process of the diffusion model.
@@
III-C Pixel Level Control with Processor
Image restoration aims to recover high-quality authentic images from low-quality images. However, real-world LQ images often suffer from complex and unknown degradation. Directly extracting features from LQ images to control the diffusion process inevitably leads to the generation of images with significant artifacts. Therefore, we propose the use of a Pixel-level Processor to mitigate the impact of degradation and obtain clear latent features from the input LQ images for controlling the diffusion process. The processor is a tiny network of four convolution layers to convert image-based conditions to 64 × 64 feature space to match the diffusion process. ControlNet [52] utilizes these latent features to generate pixel control signals in the latent domain, which are used to control the denoising process of the diffusion model. The process can be formulated as follows:
@@ III-C Pixel Level Control with Processor
where $I_{LQ}$, $F$, and $P_{i}$ represent the LQ image, preprocessing feature and pixel control respectively and $E_{processor}$, $E_{ControlNet}$ represent Pixel-level Processor and ControlNet module respectively.
What’s more, to ensure the effectiveness of the processing operation, we use RGB loss $L_{RGB}$ and FFT loss $L_{FFT}$ for supervision. Specifically, we extract intermediate features with 1/2, 1/4, and 1/8 resolutions ($F_{1/2},F_{1/4},F_{1/8}$) from the processor. Similar to PASD [30], we employ a convolutional layer to transform the feature maps at each scale into three-channel RGB space and compute the L1 distance with the corresponding high-quality images. To make further supervision and narrow gaps in the frequency domain, we make the FFT transform and calculate the loss in the frequency domain. The loss functions are computed as follows:
@@
where $I_{i}$ denotes the RGB image transformed from $F_{i}$, $I_{i}^{gt}$ denotes the ground truth at scale $i$ and $FFT$ denotes the Fast Fourier transform.
III-D Integrating Control to Denoising U-Net
Based on the above textual level and pixel level components, we have obtained text embedding $T_{text}$, image embedding $T_{image}$ and pixel control $\{P_{i}\}_{i=1}^{4}$ to guide the diffusion model’s denoising process. Unlike the original ControlNet, we introduce pixel attention to integrate pixel control into the denoising process. Additionally, text embedding continues to serve as textual prompts input to the original diffusion model’s text cross attention, leveraging the pre-trained text-to-image generation model’s prior knowledge. To enable collaboration between text embedding and image embedding, we introduce another image cross attention to integrate image embedding into the denoising U-Net. For the attention mechanism, we use $Q$, $K$, and $V$ to represent the query, key, and value features, respectively. For pixel attention, $Q$ comes from the denoising U-Net’s features, while $K$ and $V$ come from the pixel control $P_{i}$. The $K$ and $V$ of text cross attention and image cross attention come from text embedding and image embedding, respectively. The attention module’s structure is illustrated in Fig 2, with the Down Attention and Up Attention represented in green and orange, respectively. The Down Attention module includes self-attention, image cross attention, and text cross attention. The Up Attention module includes self-attention, image cross attention, text cross attention, and pixel attention.
@@ III-E Training and Inference Strategies
During the training process, the HQ image is encoded by the pre-trained VAE encoder [47] to obtain the latent code $z_{0}$. The diffusion process progressively adds noise to $z_{0}$ to generate $z_{t}$, where $t$ represents the random-sampled diffusion steps. By controlling the diffusion step $t$, the LQ image $I_{LQ}$, and the MLLM predicted prompt $T_{MLLM}$, we train the proposed image restoration network, denoted as $\epsilon_{\theta}$, to estimate the noise added to the noise layer $z_{t}$. Our optimization objective of the diffusion process is:
@@
where $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters. We set $\lambda_{1}$ and $\lambda_{2}$ equal to 0.1 and 0.01 empirically.
Specifically, the original CLIP text encoder in SD only encodes 75 tokens and cannot handle the entire generation prompts of MLLM. Therefore, we repeatedly use the text encoder to guide the encoding of all prompts. As illustrated in Fig.2, we freeze the parameters of the original SD model and only train the newly introduced modules, which include the Pixel-level Processor, ControlNet, and the newly added image attention and pixel attention within the attention module to minimize training costs.
@@ IV-A Experimental Settings
Implementation Details. We employ the pre-trained SD 1.5  [^2] https://huggingface.co/runwayml/stable-diffusion-v1-5 model as the base pre-trained T2I model. During training, we finetune the model with Adam [55] optimizer for 100K iterations. The batch size and the learning rate are set to 32 and 5e-5. All experiments are conducted on an NVIDIA A100 GPU. For inference, we adopt DDPM sampling [56] with 50 timesteps.
Training and Testing Datasets. We train our model on DIV2K [57], Flickr2K [58], OST [59], the 9K face images from FFHQ [60] and the 20K images from LSDIR [61]. We use the degradation pipeline of Real-ESRGAN [17] to synthesize LQ-HQ training pairs. To conduct a comprehensive and reliable evaluation of the real-IR task, we conduct testing using synthetic and real-world datasets. For the synthetic test dataset, we create two datasets: DIV2K-Val and LSDIR-Val, which are generated from the DIV2K validation set [57] and the LSDIR validation set [61], respectively. For each synthetic dataset, We randomly crop 2K patches from the original validation set following the same degradation pipeline utilized during training. For the real-world test datasets, we employ the two widely used datasets, RealSR [62] and DrealSR [63] for evaluation and use the same configuration as [28] to center-crop the LQ image to 128 × 128. What’s more, We also use another real-world dataset, named RealLR200 [31], which comprises 200 LQ images that have been widely used in recent works and cover a diverse range of application scenarios.
@@
Datasets Metrics BSRGAN Real-ESRGAN DASR FeMaSR LDM StableSR DiffBIR PASD SeeSR Ours
@@
Evaluation Metrics. To provide a comprehensive and effective quantitative evaluation of the different methods, we employ a range of widely used reference and non-reference metrics. In reference-based metrics, PSNR and SSIM [64] (calculated on the Y channel in YCbCr space) are fidelity metrics, while LPIPS [65] and DISTS [66] are quality evaluation metrics. FID [67] calculates the distance between the distributions of restored images and high-quality images. NIQE [68], MANIQA [69], MUSIQ [70], and CLIPIQA [71] are non-reference image quality assessment (IQA) metrics. We employ the IQA-Pytorch  [^3] https://github.com/chaofengc/IQA-PyTorch to compute these metrics.
Compared Methods. We compare our method with several state-of-the-art methods, which include GAN-based methods (BSRGAN [16], Real-ESRGAN [17], DASR [18] and FeMaSR) and diffusion-based methods (LDM [47], StableSR [28], DiffBIR [29], PASD [30] and SeeSR [31]). We use the released version of these competing methods for testing.
@@
Datasets Metrics BSRGAN Real-ESRGAN DASR FeMaSR LDM StableSR DiffBIR PASD SeeSR Ours
PSNR $\uparrow$ 24.75 24.15 25.40 23.51 23.83 23.95 23.29 24.70 23.66 23.76
SSIM $\uparrow$ 0.7401 0.7363 0.7458 0.7088 0.6857 0.7240 0.6341 0.7213 0.6952 0.7088
LPIPS $\downarrow$ 0.2656 0.2710 0.3134 0.2937 0.3159 0.2604 0.3567 0.2846 0.3004 0.2865
FID $\downarrow$ 141.26 135.18 132.62 140.02 132.67 132.10 124.81 120.97 125.07 123.21
DISTS $\downarrow$ 0.2124 0.2060 0.2202 0.2286 0.2215 0.1990 0.2298 0.2040 0.2218 0.2136
NIQE $\downarrow$ 5.6362 5.8031 6.5454 5.7685 6.5139 6.6285 5.6486 5.3805 5.3968 5.5711
MANIQA $\uparrow$ 0.3764 0.3733 0.2461 0.3632 0.3635 0.3759 0.4602 0.4413 0.5369 0.5544
MUSIQ $\uparrow$ 63.29 60.37 41.20 59.06 58.91 61.81 65.29 64.48 69.70 69.81
RealSR CLIPIQA $\uparrow$ 0.5116 0.4491 0.3200 0.5410 0.5680 0.5422 0.6558 0.5636 0.6680 0.6765
PSNR $\uparrow$ 26.39 26.28 27.24 24.85 25.57 26.70 24.74 26.29 25.90 25.31
SSIM $\uparrow$ 0.7739 0.7767 0.7995 0.7247 0.7493 0.7774 0.6155 0.7475 0.7434 0.7351
LPIPS $\downarrow$ 0.2858 0.2819 0.3099 0.3157 0.3379 0.2698 0.4348 0.3320 0.3141 0.3265
FID $\downarrow$ 155.60 147.64 155.37 157.80 155.90 151.22 164.93 154.56 146.98 152.90
DISTS $\downarrow$ 0.2144 0.2089 0.2275 0.2239 0.2256 0.2066 0.2691 0.2255 0.2298 0.2347
NIQE $\downarrow$ 6.5327 6.6932 7.5868 5.9035 7.1200 7.5441 6.0035 6.7395 6.4766 6.6904
MANIQA $\uparrow$ 0.3425 0.3436 0.2845 0.3163 0.3451 0.3188 0.4553 0.4058 0.5057 0.5255
DrealSR MUSIQ $\uparrow$ 57.17 54.27 42.41 53.71 53.73 51.36 61.46 56.12 64.73 66.27
CLIPIQA $\uparrow$ 0.5097 0.4520 0.3815 0.5642 0.5756 0.4905 0.6574 0.5682 0.6895 0.6955
NIQE $\downarrow$ 4.3656 4.1767 4.3180 4.6286 - 4.4676 3.7673 3.8752 4.2464 4.4406
MANIQA $\uparrow$ 0.3671 0.3633 0.2937 0.4083 - 0.3622 0.4734 0.4314 0.4845 0.5164
MUSIQ $\uparrow$ 64.87 62.96 55.71 64.24 - 61.36 67.27 66.96 68.37 67.72
RealLR200 CLIPIQA $\uparrow$ 0.5699 0.5409 0.4689 0.6548 - 0.5545 0.7022 0.6284 0.6550 0.6930

### v2

@@ III-B1 Analysis of the Influence of Semantic Prior
Figure 3: The comparison of different prompts and their corresponding restoration results with PASD [33]. (a) -(d) show the null prompt, BLIP prompt from LQ image, BLIP prompt from GT image and MLLM prompt predicted from LQ image and its corresponding restoration result. MLLM prompt is “A field of ripe wheat stands against a clear blue sky, creating a picturesque sight. The golden wheat glows under the sunlight, while the serene blue sky provides a beautiful backdrop.”
III-B2 Explicit-Implicit Strategy for Semantic Prior Extraction
@@
In concrete, benefiting from multimodal large language models, which possess extensive knowledge and powerful perception and reasoning capabilities, we utilize an MLLM to predict reasonable content prompts for potential blind image restoration results. Large language models can generate promising responses to user queries based on image information. We employ the pre-trained multi-modal large language model LLaVA-7B  [^1] https://huggingface.co/liuhaotian/llava-v1.5-7b to predict potential content representations from low-quality images, generating coherent and detailed textual descriptions and encoding them as text embeddings. The aforementioned process utilizes pre-trained models and does not require fine-tuning. The process can be summarized as follows:
@@
where $T_{text}$ and $I_{LQ}$ denote the text embedding and the input low-quality image respectively, $T_{input}$ denotes the text instruction to the MLLM, such as “Describe the image in a very detailed manner if we remove the degradation artifacts from the image.” and $E_{text}$ and $E_{MLLM}$ indicates the CLIP text encoder and MLLM model respectively. MLLM can take both image and text sequences as input and generate coherent descriptions as output. Compared to predicting semantic tags [34] and cognitive embeddings [35], the predicted descriptions are more detailed. In addition, the semantic prior obtained in this way is explicit and conducive to human feedback.
The inference results of the MLLM include fine-grained descriptions. To ensure that the generated image remains semantically faithful to the original low-quality image, we introduce an image embedding branch to extract visual information as a supplement. Specifically, we use CLIP image encoder to extract additional image embedding. To eliminate information about degradation artifacts present in the image embedding, we design a Refine Layer to correct the image embedding, generating image embedding that is semantically faithful to the original low-quality image but free from degradation information. The process can be expressed as follows:
@@
where $T_{image}$, $E_{image}$, and $E_{refine}$ represent the image embedding, CLIP image encoder, and the Refine Layer respectively. The Refine Layer contains three MLP layers with Layer Normalization and LeakyReLU activations (except for the final layer). The semantic prior obtained in this way is implicit and can better reflect the semantic information in the input image that cannot be represented by explicit text.
Ultimately, the inferred explicit text embedding ($T_{text}$) and implicit image embedding ($T_{image}$) collaborate as the semantic prior to control the restoration process of the diffusion model.
GT LQ BSRGAN[16] Real-ESRGAN [17] StableSR[31] DiffBIR[32] PASD[33] SeeSR[34] Ours
@@
III-C Structure Prior Extraction with Constrained Processor
Image restoration aims to recover high-quality authentic images from low-quality images. Different from general generation tasks, it requires the generated image to be structurally faithful to the low-quality image. However, due to real-world LQ images often suffer from complex and unknown degradation, directly extracting features from LQ images to control the diffusion process inevitably leads to the generation of images with undesirable artifacts. Therefore, we propose a multi-level constrained Processor to mitigate the impact of degradation and obtain degradation-independent latent structure features from the input LQ images for controlling the diffusion process. The Processor is a tiny network of four convolution layers to convert image-based conditions to 64 × 64 feature space to match the diffusion process. ControlNet [56] utilizes these latent features to generate structure control signals in the latent domain, which are used to control the denoising process of the diffusion model. The process can be formulated as follows:
@@ III-C Structure Prior Extraction with Constrained Processor
where $I_{LQ}$, $F$, and $P_{i}$ represent the LQ image, preprocessing feature and structure control respectively, and $E_{processor}$, $E_{ControlNet}$ represent the Processor and ControlNet module respectively.
What’s more, to ensure the effectiveness of the processing operation, we use the multi-level RGB loss $\mathcal{L}_{RGB}$ and FFT loss $\mathcal{L}_{FFT}$ for supervision. Specifically, we extract intermediate features with 1/2, 1/4, and 1/8 resolutions ($F_{1/2},F_{1/4},F_{1/8}$) from the processor. Similar to PASD [33], we employ a convolutional layer to transform the feature maps at each scale into three-channel RGB space and compute the L1 distance with the corresponding high-quality images. To make further supervision and narrow gaps in the frequency domain, we make the FFT transform and calculate the loss in the frequency domain. The loss functions are computed as follows:
@@
where $I_{i}$ denotes the RGB image transformed from $F_{i}$, $I_{i}^{gt}$ denotes the ground truth at scale $i$ and $FFT$ denotes the Fast Fourier transform. In summary, $\mathcal{L}_{RGB}$ and $\mathcal{L}_{FFT}$ implement constraints on different scales and dimensions of the processor, which helps to extract degradation-independent structure features.
III-D Prior-Guided Attention Module
Based on the above semantic and structure priors extraction modules, we have obtained text embedding $T_{text}$, image embedding $T_{image}$, and structure control $\{P_{i}\}_{i=1}^{4}$. All the prior information will be used to guide the diffusion model’s denoising process to generate the desired restoration results.
We design a Prior-Guided Attention module to integrate the priors. The Prior-Guided Attention module is shown at the bottom of Fig 2 and has a little difference in the Down Attention and Up Attention. The Down Attention represented in green includes self-attention, image cross attention, and text cross attention while the Up Attention represented in orange includes self-attention, image cross attention, text cross attention, and structure attention. Following the setting of the ControlNet, we only introduce the structure attention in the Up Attention of the denoising U-Net. Additionally, different from the original ControlNet which only introduces structure control by addition, we introduce structure attention to integrate structure control into the denoising process. This helps the generated images to be more faithful to the input images [33]. Text embedding is integrated into the diffusion process through text cross attention, leveraging the pre-trained text-to-image generation model’s prior knowledge. To enable collaboration between text embedding and image embedding, we introduce another image cross attention to integrate image embedding into the denoising U-Net.
For the attention mechanism, we use $Q$, $K$, and $V$ to represent the query, key, and value features, respectively. For structure attention, $Q$ comes from the denoising U-Net’s features, while $K$ and $V$ come from the structure control $P_{i}$. The $K$ and $V$ of text cross attention and image cross attention come from text embedding and image embedding, respectively. The Prior-Guided Attention module effectively integrates the semantic and structure priors to the diffusion model to generate HQ images.
LQ BSRGAN[16] Real-ESRGAN[17] StableSR[31] DiffBIR[32] PASD[33] SeeSR[34] SUPIR[42] Ours
@@ III-E Training and Inference Strategies
During the training process, the HQ image is encoded by the pre-trained VAE encoder [52] to obtain the latent code $z_{0}$. The diffusion process progressively adds noise to $z_{0}$ to generate $z_{t}$, where $t$ represents the random-sampled diffusion steps. By controlling the diffusion step $t$, the LQ image $I_{LQ}$, and the MLLM predicted prompt $T_{MLLM}$, we train the proposed diffusion-based restoration network, denoted as $\epsilon_{\theta}$, to estimate the noise added to the noise layer $z_{t}$. Our optimization objective of the diffusion process is:
@@
where $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters.
Specifically, the original CLIP text encoder in SD only encodes 75 tokens and cannot handle the entire generation prompts of MLLM. Therefore, we repeatedly use the text encoder to guide the encoding of all prompts. As illustrated in Fig.2, we freeze the parameters of the original SD model and only train the newly introduced modules, which include the Constrained Processor, ControlNet, and the newly added image attention and structure attention within the attention module to minimize training costs.
@@ IV-A Experimental Settings
Implementation Details. We employ the pre-trained SD 1.5  [^2] https://huggingface.co/runwayml/stable-diffusion-v1-5 model as the base pre-trained T2I model. During training, we finetune the model with Adam [58] optimizer for 100K iterations. The batch size and the learning rate are set to 32 and 5e-5. For the hyperparameters in the loss function, we set $\lambda_{1}$ and $\lambda_{2}$ equal to 0.1 and 0.01 empirically. All experiments are conducted on an NVIDIA A100 GPU. For inference, we adopt DDPM sampling [59] with 50 timesteps.
Training and Testing Datasets. We train our model on DIV2K [60], Flickr2K [61], OST [62], the 9K face images from FFHQ [63] and the 20K images from LSDIR [64]. We use the degradation pipeline of Real-ESRGAN [17] to synthesize LQ-HQ training pairs. The pipeline consists of a second-order degradation process and each process randomly adds blur, resize, noise, and JPEG compression artifacts. The blur degradation involves isotropic Gaussian and anisotropic Gaussian kernels. The resize degradation involves area resize, bilinear interpolation and bicubic resize. The noise degradation involves additive Gaussian noise and Poisson noise. To conduct a comprehensive and reliable evaluation of the real-IR task, we conduct testing using synthetic and real-world datasets. For the synthetic test dataset, we create two datasets: DIV2K-Val and LSDIR-Val, which are generated from the DIV2K validation set [60] and the LSDIR validation set [64], respectively. For each synthetic dataset, We randomly crop 2K patches from the original validation set following the same degradation pipeline utilized during training. For the real-world test datasets, we employ the two widely used datasets, RealSR [23] and DrealSR [65] for evaluation and use the same configuration as [31] to center-crop the LQ image to 128 × 128. What’s more, We also use another real-world dataset, named RealLR200 [34], which comprises 200 LQ images that have been widely used in recent works and cover a diverse range of application scenarios.
@@
Datasets Metrics BSRGAN[16] Real-ESRGAN[17] DASR[18] FeMaSR[19] LDM[52] StableSR[31] DiffBIR[32] PASD[33] SeeSR[34] Ours
@@
Evaluation Metrics. To provide a comprehensive and effective quantitative evaluation of the different methods, we employ a range of widely used reference and non-reference metrics, including PSNR, SSIM[66] (calculated on the Y channel in YCbCr space), LPIPS[67], FID[68], DISTS[69], CLIP-I[70], NIQE[71], MANIQA[72], MUSIQ[73], and CLIPIQA[74]. For reference metrics, PSNR and SSIM are used to assess the similarity between generated images and ground truth images, while LPIPS and DISTS are employed to evaluate the perceptual quality of the generated images. CLIP-I estimates semantic accuracy by evaluating CLIP scores between generated images and ground truth images. For non-reference metrics, NIQE assesses the naturalness of images by utilizing statistical features derived from natural images. MANIQA evaluates the quality of images from both global and local perspectives using a multi-dimensional attention network. MUSIQ captures multi-scale image representations to provide a comprehensive assessment. CLIPIQA leverages the pre-trained vision-language model CLIP to evaluate image perceptual quality based on textual descriptions and visual features. These metrics offer objective measurements to evaluate the performance of different methods in terms of image similarity and perceptual quality, enabling a comprehensive assessment of the generated images. Non-reference metrics, in particular, are capable of better measuring the perceptual quality of images. We employ the IQA-Pytorch  [^3] https://github.com/chaofengc/IQA-PyTorch to compute these metrics.
@@
Datasets Metrics BSRGAN[16] Real-ESRGAN[17] DASR[18] FeMaSR[19] LDM[52] StableSR[31] DiffBIR[32] PASD[33] SeeSR[34] SUPIR[42] Ours
PSNR $\uparrow$ 24.75 24.15 25.40 23.51 23.83 23.95 23.29 24.70 23.66 23.08 23.76
SSIM $\uparrow$ 0.7401 0.7363 0.7458 0.7088 0.6857 0.7240 0.6341 0.7213 0.6952 0.6470 0.7088
LPIPS $\downarrow$ 0.2656 0.2710 0.3134 0.2937 0.3159 0.2604 0.3567 0.2846 0.3004 0.3585 0.2865
FID $\downarrow$ 141.26 135.18 132.62 140.02 132.67 132.10 124.81 120.97 125.07 128.45 123.21
DISTS $\downarrow$ 0.2124 0.2060 0.2202 0.2286 0.2215 0.1990 0.2298 0.2040 0.2218 0.2414 0.2136
CLIP-I $\uparrow$ 0.8635 0.8756 0.8768 0.8759 0.8681 0.8838 0.8770 0.8895 0.8847 0.8703 0.8956
NIQE $\downarrow$ 5.6362 5.8031 6.5454 5.7685 6.5139 6.6285 5.6486 5.3805 5.3968 6.77 5.5711
MANIQA $\uparrow$ 0.3764 0.3733 0.2461 0.3632 0.3635 0.3759 0.4602 0.4413 0.5369 0.5041 0.5544
MUSIQ $\uparrow$ 63.29 60.37 41.20 59.06 58.91 61.81 65.29 64.48 69.70 62.21 69.81
RealSR CLIPIQA $\uparrow$ 0.5116 0.4491 0.3200 0.5410 0.5680 0.5422 0.6558 0.5636 0.6680 6.6573 0.6765
PSNR $\uparrow$ 26.39 26.28 27.24 24.85 25.57 26.70 24.74 26.29 25.90 24.60 25.31
SSIM $\uparrow$ 0.7739 0.7767 0.7995 0.7247 0.7493 0.7774 0.6155 0.7475 0.7434 0.6487 0.7351
LPIPS $\downarrow$ 0.2858 0.2819 0.3099 0.3157 0.3379 0.2698 0.4348 0.3320 0.3141 0.4137 0.3265
FID $\downarrow$ 155.60 147.64 155.37 157.80 155.90 151.22 164.93 154.56 146.98 152.19 152.90
DISTS $\downarrow$ 0.2144 0.2089 0.2275 0.2239 0.2256 0.2066 0.2691 0.2255 0.2298 0.2644 0.2347
CLIP-I $\uparrow$ 0.8807 0.8935 0.8924 0.8912 0.8787 0.8921 0.8843 0.8837 0.8851 0.8854 0.8960
NIQE $\downarrow$ 6.5327 6.6932 7.5868 5.9035 7.1200 7.5441 6.0035 6.7395 6.4766 7.69 6.6904
MANIQA $\uparrow$ 0.3425 0.3436 0.2845 0.3163 0.3451 0.3188 0.4553 0.4058 0.5057 0.4663 0.5255
DrealSR MUSIQ $\uparrow$ 57.17 54.27 42.41 53.71 53.73 51.36 61.46 56.12 64.73 57.55 66.27
CLIPIQA $\uparrow$ 0.5097 0.4520 0.3815 0.5642 0.5756 0.4905 0.6574 0.5682 0.6895 0.6469 0.6955
NIQE $\downarrow$ 4.3656 4.1767 4.3180 4.6286 - 4.4676 3.7673 3.8752 4.2464 4.3250 4.4406
MANIQA $\uparrow$ 0.3671 0.3633 0.2937 0.4083 - 0.3622 0.4734 0.4314 0.4845 0.4246 0.5164
MUSIQ $\uparrow$ 64.87 62.96 55.71 64.24 - 61.36 67.27 66.96 68.37 63.13 67.72
RealLR200 CLIPIQA $\uparrow$ 0.5699 0.5409 0.4689 0.6548 - 0.5545 0.7022 0.6284 0.6550 0.6114 0.6930
Compared Methods. We compare our method with several state-of-the-art methods, which include GAN-based methods (BSRGAN [16], Real-ESRGAN [17], DASR [18] and FeMaSR) and diffusion-based methods (LDM [52], StableSR [31], DiffBIR [32], PASD [33] and SeeSR [34]). Additionally, We also make a comparison with the MLLM assisted method SUPIR on the three real-world datasets. We use the released version of these competing methods for testing.